# Quantised Global Autoencoder: A Holistic Approach to Representing Visual Data

### v1

@@ Abstract
In quantised autoencoders, images are usually split into local patches, each encoded by one token. This representation is redundant in the sense that the same number of tokens is spend per region, regardless of the visual information content in that region. Adaptive discretisation schemes like quadtrees are applied to allocate tokens for patches with varying sizes, but this just varies the region of influence for a token which nevertheless remains a local descriptor. Modern architectures add an attention mechanism to the autoencoder which infuses some degree of global information into the local tokens. Despite the global context, tokens are still associated with a local image region. In contrast, our method is inspired by spectral decompositions which transform an input signal into a superposition of global frequencies. Taking the data-driven perspective, we learn custom basis functions corresponding to the codebook entries in our VQ-VAE setup. Furthermore, a decoder combines these basis functions in a non-linear fashion, going beyond the simple linear superposition of spectral decompositions.
We can achieve this global description with an efficient transpose operation between features and channels and demonstrate our performance on compression.
@@ 1 Introduction
The standard convolution operation on images only takes a local neighbourhood of pixels into account. A vanilla convolutional autoencoder therefore produces an encoding that stores aggregated local information in each latent vector, as larger receptive fields are costly. With the lack of global information, each part of an image needs the same amount of encodings to describe it properly. Since images often have parts that contain much information, while other parts do not have a high information content, encoding each part of the image with the same amount of tokens is counter intuitive: Some tokens might contain little information, while other regions might not be properly represented because of a lack of tokens in the vicinity.
@@
Instead of encoding patches with often vastly different amounts of information content using the same amounts of tokens, our global approach on compression shows better performance, even outperforming methods that learn adaptive hierarchical descriptions[12]. We demonstrate autoregressive generation as a downstream task.
@@
As we introduce global elements into the VQ-VAE framework, we call our approach Quantised Global Variational Autoencoder (QG-VAE).
@@ 2 Related Work
We discuss Traditional Compression Methods for image compression and learned methods based around Autoencoders, as we combine ideas from both domains. We further elaborate the different approaches in learned autoencoder-based approaches, then dive into Quantised Methods that try to increase the information content of the representation, as we aim to produce a quantised and finite representation with our approach.
@@ 2.1 Traditional Compression Methods
In traditional image compression, images are often decomposed into different base frequency bands, then expressed as a linear combination. Exemplary, Watson et al. use the discrete cosine transformation[35] that was later used for the JPEG compression standard[34]. Similarly, the Laplacian Pyramid[3] expresses an image as a set of frequency images ordered from coarse to fine, finding application e.g. in texture synthesis [10]. As a related approach, Eigenfaces[30] compute a basis set of faces, then express faces as a linear combination, only storing coefficients of the basis functions.
While these methods are fast and do not require training a neural network, they all only express images as a linear combination of components, never in a context sensitive and holistic manner.
@@ 3.1 Background
The core idea of this transform is to represent a (spatial) signal as a superposition of frequency functions. If we restrict the signal to be discrete (and periodic), the frequency spectrum is also discrete (and periodic) leading to the so-called Discrete Fourier Transform (DFT):
@@
where $I_{x,y}$ describes a pixel of an image $I$ that has the size $W\times H$. By using (complex) sine and cosine functions as basis functions of a linear function space, each input signal can be effectively approximated by a linear combination of frequency basis functions. The coefficients of this linear combination are the frequency components (complex amplitudes). If we reconstruct the signal from only a sub sequence of the frequency components, we obtain a low frequency approximation of the input. More generally, we can express the frequency decomposition $F$ for $n$ frequencies as a set of independent frequency band computations with:
@@
We call the resulting factors of our decomposition pseudo frequency tokens, as they differ to a regular frequency decomposition in two main aspects. While in the Fourier transform, the coefficients are computed by evaluating a dot product between the input and a fixed frequency basis functions, we replaces this by a learned neural network and quantise these coefficients. While learning an individual function per pseudo frequency would be ideal, this would, for $C$ frequency bands, require learning $C$ different global functions that process the input image. This is impractical for our purposes due to a high number of parameters and hence difficulties in generalising across multiple images. It also does not allow processing the different pseudo frequencies in context to each other. Instead, we resort to a small change that makes our approach feasible. Where DFT applies a different function to the same image through the changed exponent to produce each frequency, we instead use the same (learned) function $K$ at our core to produce the different pseudo frequencies. We obtain these pseudo frequencies $S_{i},i\in{0,1,...,C}$ by applying $K$ to $C$ different feature maps of the input image, with $E$ producing a feature map of the input:
@@ 3.2 Quantised Global Autoencoder
The following paragraphs describe these steps in more detail.
@@
To ensure we have all the information distributed across the feature maps that each form their own pseudo frequency, we apply a U-Net [24] before and after the switch from local to global features, referred to as Encoder $E$ and Decoder $D$. These U-Nets, together with the affine transformation $K$, then learn what corresponds to the projection to the basis functions. To keep memory requirements reasonable, we omit some of the upscaling (encoder) or downscaling (decoder) steps for larger images, e.g. turning our $[B\times 3\times W\times H]$ dimensional input into a $[B\times C\times\frac{W}{2^{f}}\times\frac{H}{2^{f}}]$ feature map for a downscaling factor $f$.
@@
To then quantise, we use the same approach as traditional VQ-VAEs [31]: Rounding to the nearest codeword from our codebook, and then both regularising so that the encoder produces an output close to the codewords ($L_{comm}$) and so that the codewords move closer to what the encoder produces ($L_{code}$). We use the same loss formulation that uses the distance function to nearest codeword $d$ from the codebook $e$:
@@ 3.4 Obtaining a Meaningful Decomposition of the Latent Space
While our approach already provides a decomposition into pseudo frequencies that is somewhat structured, we can order this from important to unimportant. This extra structure can be an additional help to understand our approach and is yet again inspired by classic frequency decomposition.
@@
4 Evaluation
We run all our benchmarks on a single Nvidia GeForce RTX 2080 Ti, demonstrating that our approach requires no expensive hardware or excessive amounts of VRAM. We further use a U-Net with residual blocks from [26] and downscaling/upsampling instead of max pooling/transposed convolution to make our code less intricate. We include an exemplary one-click Jupyter Notebook with the right settings [^1] Code can be found at https://www.graphics.rwth-aachen.de/media/resource_files/demo.ipynb.
4.1 Benchmarks in Compression
We interpret these results in the context of global versus local descriptions: An image is no longer composed of patches that possibly have visible seams that can be picked up by the NN-based metrics like FID and LPIPS. Our approach is biased towards learning an accurate, holistic representation through its global feature maps instead of local patches that form a token. We also observe that features in produced images compared to a regular VQ-VAE might still not be correct, i.e. offset or wrong, but they look much sharper than at similar PSNR values. This heavily shows in the FID scores of CIFAR-10. For reference, consider high-frequency details like the legs of the horse rider in Fig. 4. Also note that to compute the FID values, the latent space of a model trained on ImageNet is used, which does not contain close up features of faces, and should hence be taken with a grain of salt on the CelebA benchmark. We provide extra qualitative results in Sec. G, in particular on larger datasets, resolutions, and number of tokens. We also do some qualitative analysis and compare to a VQ-VAE directly to better show the visual improvements of our approach.
@@ 4.1.1 Comparison to Adaptive Refinement
4.2 Applications of our Latent Space
While compression is our main focus, we did some basic exploration of using our tokenised $1$D latent space for generation. We theorise that our natively linear latent code is beneficial for generation: Autoregressively predicting token by token, and hence, patch by patch of in a $2$D grid, does not come native for a model designed for 1D sequences. We provide some (early) results on generating CIFAR images with an autoregressive transformer[13], hinting at the potential for future work, in Sec. H.
@@ 4.2 Applications of our Latent Space
Figure 7: Learned feature maps of our approach: We normalised the individual $C$ many $\frac{W}{2^{f}}$-by-$\frac{H}{2^{f}}$ feature maps from the step before they would be compressed into pseudo frequency tokens by the affine transformation, then visualised $16$ randomly chosen ones here, all from the same image as Fig. 6, but not ordered, i.e. without Sec. 3.4.
@@ 4.3 Interpretation of our Holistic Representation
We generally believe our better compression rate to be rooted in holistic instead of local descriptions of an input. We explicitly demonstrate how our learned latent space produces global features instead of localised ones with additional experiments, from which we can draw further conclusions about the properties of our learned space.
First, we visualise the feature maps, i.e. the $C$ many different channels of our $[B\times C\times\frac{W}{2^{f}}\times\frac{H}{2^{f}}]$ tensor before compressing each $C$ dimension into a token by affine transformation and quantisation in Fig. 7. This shows how each token captures different global properties, following the a holistic strategy to not just encode an input through local patches.
Second, we average the change for all possible replacements for a token in the output in Fig. 1. We observe very global changes for single token changes, compared to very local changes for VQ-VAE. Similarly, we can slightly change an input image to provoke large changes in the learned latent space: For a simple change of $20$-by-$20$ pixels in an $128$-by-$128$ pixel image, we observe a change of $27$ tokens, while we have $\frac{128^{2}}{256}=64$ pixels per token. An example can be found in Sec. F.
Third, we train an ordered latent space according to Sec. 3.4, with the first $k$ tokens always holding the largest amount of content in the least squares sense. Then, we show the learned decomposition in Fig. 6, showing how gradually the image changes from an average to the target image.
Notably, even with this enforced decomposition, our approach has properties that are not typical for a simple linear combination. As an example, we consider e.g. Eigenfaces[30] that compute an ordered set of face coefficients from high to low importance via PCA. Hence, adding new coefficients of base elements provides additional detail, while our approach is able to completely change attributes instead (e.g. gradually moving the smile with more tokens in Fig. 6). However, as our ordering regularisation means that each first $k$ tokens of pseudo frequency bands must also be interpretable on their own, we observe some decay in reconstruction quality compared to our holistic version without regularisation.
@@ 4.4 Limitations, Discussion, Future Work
Further, our approach does generally require more parameters, as two U-Nets as backbone are more expensive as e.g. convolution and pooling of a regular VQ-VAE. However, especially for smaller resolutions (e.g. CIFAR), we find that our approach performs similarly with the same number of parameters. Even with those higher numbers of parameters, we find that a U-Net with at most $128$ channels is not extra ordinarily expensive given the shown performance, in particular considering that our vanilla loss function identical to a regular VQ-VAE[31] is much cheaper to compute than e.g. the mathematical regularisation of [33]. We show the exact parameter combinations in Sec. A and discuss details in Sec. B. Additionally, we do not optimally utilise our codebooks: We do not ensure that our codebook resets are optimal, as we do not ensure that we split a codebook entry in an optimal way (Fig. 3 shows an optimal split). This can lead to potentially having multiple codewords that are very similar, but are still each with a high gradient. We only reset unused codebook entries instead of resetting very rarely used ones. Also, we decided not to include attention[1], as to not dilute that the architecture itself is able to produce a global representation. We leave this for future work to keep our baseline approach simple.
@@ 5 Conclusion
We propose a holistic approach to autoencoding, capturing global features in quantised tokens. Opposing to traditional decompositions, we process these pseudo frequency tokens together. We believe that this holistic take on autoencoding has the potential to spawn a whole new branch of autoencoders, as we demonstrate success in multiple benchmarks with a rather simple formulation. We do so without any powerful mathematical regularisations like Wasserstein loss or recent architectural improvements like attention or artificial sharpeners.
@@ A Benchmark Parameters
Default ImageNet128128128128[5] $256$ $32$-by-$32$ / $f=2$ $64$ $2$ $128$ $1000$ $100$

### v2

@@ Abstract
Quantised autoencoders usually split images into local patches, each encoded by one token. This representation is redundant because the same number of tokens is spent per region, regardless of the visual information content in that region. Adaptive discretisation schemes like quadtrees allocate tokens for patches with varying sizes, but this just varies the region of influence for a token which nevertheless remains a local descriptor. Modern architectures add an attention mechanism to the autoencoder to infuse some degree of global information into the local tokens. Despite the global context, tokens are still associated with a local image region. In contrast, our method is inspired by spectral decompositions which transform an input signal into a superposition of global frequencies. Taking the data-driven perspective, we learn custom basis functions corresponding to the codebook entries in our VQ-VAE setup. Furthermore, a decoder non-linearly combines these basis functions, going beyond the simple linear superposition of spectral decompositions.
We achieve this global description with an efficient transpose operation between features and channels and demonstrate our performance on compression. We further show that our space can help to significantly improve generation.
@@ 1 Introduction
The standard convolution operation on images takes only a local neighbourhood of pixels into account. A vanilla convolutional autoencoder therefore produces an encoding that stores aggregated local information in each latent vector, as enlarging receptive fields becomes costly. Without global information, each part of an image needs the same amount of encodings to describe it properly. Since images often have varying amounts of information content per region, encoding each part of the image with the same amount of tokens is counter intuitive. Some tokens might contain little information, while other regions might not be properly represented because of a lack of tokens in the vicinity.
@@
  • We demonstrate how our approach can boost performance in downstream tasks like generation, and how it is compatible with autoencoder improvements like sharpening.
@@
Instead of encoding patches with often vastly different amounts of information content using the same amounts of tokens, our global approach on compression shows better performance, even outperforming methods that learn adaptive hierarchical descriptions[12]. We demonstrate autoregressive generation as a downstream task that greatly benefits from our latent space.
@@
As we introduce global elements into the VQ-VAE framework, we call our approach Quantised Global Variational Autoencoder (QG-VAE). When adding additional sharpening in the spirit of VQGAN[7], we refer to the approach as QGGAN.
@@ 2 Related Work
We discuss Traditional Compression Methods for image compression and learned methods based around Autoencoders, as we combine ideas from both domains. We further elaborate the different learned autoencoder-based approaches, then dive into Quantised Methods that try to increase the information content of the representation, as we aim to produce a quantised and finite representation.
@@ 2.1 Traditional Compression Methods
In traditional image compression, images are often decomposed into different base frequency bands, then expressed as a linear combination thereof. Exemplary, Watson et al. use the discrete cosine transformation[35] that was later used for the JPEG compression standard[34]. Similarly, the Laplacian Pyramid[3] expresses an image as a set of frequency images ordered from coarse to fine, finding application e.g. in texture synthesis [10]. As a related approach, Eigenfaces[30] compute a basis set of faces, then express faces as a linear combination, only storing coefficients of the basis functions.
These methods are fast and do not require training a neural network, but all only express images as a linear combination of components, never in a context sensitive and holistic manner.
@@ 3.1 Background
The idea of this transform is to represent a (spatial) signal as a superposition of frequency functions. If we restrict the signal to be discrete (and periodic), the frequency spectrum is also discrete (and periodic) leading to the so-called Discrete Fourier Transform (DFT):
@@
where $I_{x,y}$ describes a pixel of an image $I$ of the size $W\times H$. By using (complex) sine and cosine functions as basis functions of a linear function space, each input signal can be effectively approximated by a linear combination of frequency basis functions. The coefficients of this linear combination are the frequency components (complex amplitudes). If we reconstruct the signal from only a sub sequence of the frequency components, we obtain a low frequency approximation of the input. More generally, we can express the frequency decomposition $F$ for $n$ frequencies as a set of independent frequency band computations with:
@@
We call the resulting factors of our decomposition pseudo frequency tokens, as they differ to a regular frequency decomposition in two main aspects. While in the Fourier transform, the coefficients are computed by evaluating a dot product between the input and a fixed frequency basis functions, we replace this by a learned neural network and quantise these coefficients. While learning an individual function per pseudo frequency would be ideal, this would, for $C$ frequency bands, require learning $C$ different global functions that process the input image. This is impractical because a high number of parameters and hence difficulties in generalising across multiple images. It also does not allow processing the different pseudo frequencies in context to each other. Instead, we resort to a small change that makes our approach feasible. Where DFT applies a different function to the same image through the changed exponent to produce each frequency, we instead use the same (learned) function $K$ at our core to produce the different pseudo frequencies. We obtain these pseudo frequencies $S_{i},i\in{0,1,...,C}$ by applying $K$ to $C$ different feature maps of the input image, with $E$ producing a feature map of the input:
@@ 3.2 Quantised Global Autoencoder
We now describe these steps in more detail.
@@
To ensure we distribute all the information across the feature maps that each form their own pseudo frequency, we apply a U-Net [24] before and after the switch from local to global features, referred to as Encoder $E$ and Decoder $D$. These U-Nets, together with the affine transformation $K$, learn what corresponds to the projection to the basis functions. To keep memory requirements reasonable, we omit some of the upscaling (encoder) or downscaling (decoder) steps for larger images, e.g. turning our $[B\times 3\times W\times H]$ dimensional input into a $[B\times C\times\frac{W}{2^{f}}\times\frac{H}{2^{f}}]$ feature map for a downscaling factor $f$.
@@
To quantise, we use the same approach as traditional VQ-VAEs [31]: Rounding to the nearest codeword from our codebook, and then both regularising so that the encoder produces an output close to the codewords ($L_{comm}$) and so that the codewords move closer to what the encoder produces ($L_{code}$). We use the same loss formulation that uses the distance function to nearest codeword $d$ from the codebook $e$:
@@ 3.4 Obtaining a Meaningful Decomposition of the Latent Space
While our approach already provides a decomposition into pseudo frequencies that is somewhat structured, we can order this from important to unimportant. This extra structure can be an additional help to understand our representation and is yet again inspired by classic frequency decomposition.
@@
Input QG-VAE (ours) QGGAN (ours)
Figure 5: Examples on ImageNet $128$ with sharpening, from left to right: Input, output (QG-VAE), output (QGGAN). Each image has only $256$ tokens ($288$ bytes).
4 Evaluation
We run all our benchmarks on a single Nvidia GeForce RTX 2080 Ti, demonstrating that our approach requires no expensive hardware or excessive amounts of VRAM. We further use a U-Net with residual blocks from [26] and downscaling/upsampling instead of max pooling/transposed convolution to make our code less intricate. We provide our code as simple, one-clickable Jupyter Notebooks on GitHub [^1] https://github.com/DaiDaiLoh/QG-VAE.
4.1 Benchmarks in Compression
We interpret these results in the context of global versus local descriptions: An image is no longer composed of patches that possibly have visible seams that can be picked up by the NN-based metrics like FID and LPIPS. Our approach is biased towards learning an accurate, holistic representation through its global feature maps instead of local patches that form a token. We also observe that features in produced images compared to a regular VQ-VAE might still not be correct, i.e. offset or wrong, but they look much sharper than at similar PSNR values. This heavily shows in the FID scores of CIFAR-10. For reference, consider high-frequency details like the legs of the horse rider in Fig. 4. Also note that to compute the FID values, the latent space of a model trained on ImageNet is used, which does not contain close up features of faces, and should hence be taken with a grain of salt on the CelebA benchmark. We provide extra qualitative results in Sec. G, in particular on larger datasets, resolutions, and number of tokens. We also do some qualitative analysis and compare to a VQ-VAE directly to better show the visual improvements of our QG-VAE.
@@ 4.1.1 Comparison to Adaptive Refinement
4.2 Useability of our Latent Space
To show that our autoencoder framework is both compatible to common extensions and useful for downstream tasks, we demonstrate sharpening and autoregressive generation.
Sharpening
For image compression, high-frequency details are often lost in compression, e.g. the scales of a fish are not reconstructed properly. While reconstructing these high-frequency details is not feasible, approaches like VQGAN[7] produce new high frequency details by adding an additional discriminator to sharpen the output, possibly inventing similar details like the ones lost in compression. To demonstrate that our approach is just as flexible as a regular autoencoder when it comes to compatibility with other approaches, we train a VQGAN[7]-inspired approach that uses our autoencoder as its backbone, which we further refer to as QGGAN. Examples can be found in Fig. 5 and Fig. 15. We report an FID of $11.42$ for our QGGAN and $57.16$ for our trained VQGAN on ImageNet128 with only $256$ tokens after training for $15$ epochs, both without attention. We attribute the difference to more global features, that not only increase compression, but also make the recognition of global context easier to refine small details (e.g. refining the surface of a cat as fur once the network recognises the cat).
Figure 6: Generation quality when training the same transformer on our pseudo frequency tokens of our QG-VAE on CIFAR-10[15] versus training on tokens produced by a VQ-VAE[31]. Our approach performs better, producing lower FID. We compute on $10,000$ unconditional samples each.
Generation
[Uncaptioned image] While compression is our main focus, we did some basic exploration of using our tokenised $1$D latent space for generation (examples in the inset). We measure that our natively linear latent code is beneficial for generation in Fig. 6: Autoregressively predicting token by token, and hence, patch by patch of in a $2$D grid, does not come native for a model designed for 1D sequences. We provide results on generating CIFAR images with an autoregressive transformer[13], hinting at the potential for future work. After an initial delay, which we attribute to more complex relations in global instead of local tokens, e.g. predicting a blue sky token next to a blue sky token is initially trivial, the same transformer on tokens of our QG-VAE approach vastly outperforms the VQ-VAE tokens. We show additional examples and explanations in Sec. H.
@@ 4.2 Useability of our Latent Space
Figure 9: Learned feature maps of our approach: We normalised the individual $C$ many $\frac{W}{2^{f}}$-by-$\frac{H}{2^{f}}$ feature maps at the layer before they would be compressed into pseudo frequency tokens by the affine transformation. We then visualised $16$ randomly chosen ones here, all from the same image as Fig. 8, but not ordered, i.e. without Sec. 3.4.
@@ 4.3 Interpretation of our Holistic Representation
We generally believe our better compression rate to be rooted in holistic instead of local descriptions of an input. With additional experiments, we demonstrate how our learned latent space produces global features instead of localised ones with additional experiments, from which we can draw further conclusions about the properties of our learned space.
First, we visualise the feature maps, i.e. the $C$ many different channels of our $[B\times C\times\frac{W}{2^{f}}\times\frac{H}{2^{f}}]$ tensor before compressing each $C$ dimension into a token by affine transformation and quantisation in Fig. 9. This shows that each token captures different global properties, following the holistic strategy to not just encode an input through local patches.
Second, we average the change for all possible replacements for a token in the output in Fig. 1. We observe very global changes for single token changes, compared to very local changes for VQ-VAE. Similarly, we can slightly change an input image to cause large changes in the learned latent space: For a simple change of $20$-by-$20$ pixels in an $128$-by-$128$ pixel image, we observe a change of $27$ tokens, while we have $\frac{128^{2}}{256}=64$ pixels per token. An example can be found in Sec. F.
Third, we train an ordered latent space according to Sec. 3.4, with the first $k$ tokens always holding the largest amount of content in the least squares sense. Then, we show the learned decomposition in Fig. 8. This demonstrates how gradually the image changes from an average to the target image.
Notably, even with this enforced decomposition, our approach has properties that are not typical for a simple linear combination. As an example, we consider e.g. Eigenfaces[30] that compute an ordered set of face coefficients from high to low importance via PCA. Hence, adding new coefficients of base elements provides additional detail, while our approach is able to completely change attributes instead (e.g. gradually moving the smile with more tokens in Fig. 8). However, our ordering regularisation means that each first $k$ tokens of pseudo frequency bands must also be interpretable on their own. From this, we observe some decay in reconstruction quality compared to our holistic version without regularisation.
@@ 4.4 Limitations, Discussion, Future Work
Further, QG-VAE does generally require more parameters, as two U-Nets as backbone are more expensive as e.g. convolution and pooling of a regular VQ-VAE. However, especially for smaller resolutions (e.g. CIFAR), we find that our approach performs similarly with the same number of parameters. Even with those higher numbers of parameters, we find that a U-Net with at most $128$ channels is not extra ordinarily expensive given the shown performance, in particular considering that our vanilla loss function identical to a regular VQ-VAE[31] is much cheaper to compute than e.g. the mathematical regularisation of [33]. We show the exact parameter combinations in Sec. A and discuss details in Sec. B. Additionally, we do not optimally use our codebooks: We do not ensure that our codebook resets are optimal, as we do not ensure that we split a codebook entry in an optimal way (Fig. 3 shows an optimal split). This can lead to potentially having multiple codewords that are very similar, but are still each with a high gradient. We only reset unused codebook entries instead of resetting very rarely used ones. Also, we decided not to include attention[1], as to not dilute that the architecture itself is able to produce a global representation. We leave this for future work to keep our baseline approach simple.
@@ 5 Conclusion
We propose a holistic approach to autoencoding, capturing global features in quantised tokens. Opposing to traditional decompositions, we process these pseudo frequency tokens together. We believe that this holistic take on autoencoding has the potential to spawn a whole new branch of autoencoders, as we demonstrate success in multiple benchmarks with a rather simple formulation. We do so without any powerful mathematical regularisations like Wasserstein loss or recent architectural improvements like attention. We further demonstrate that our approach improves generation significantly, and that it is compatible to autoencoder improvements like additional sharpening.
@@ A Benchmark Parameters
Default ImageNet128128128128[5] $256$ $32$-by-$32$ / $f=2$ $64$ $2$ $128$ $1000$ $15$
QGGAN ImageNet128128128128[5] $256$ $16$-by-$16$ / $f=3$ $128$ $2$ $128$ $1000$ $15$